A rectangular loop of wire is placed next to a str
Answer #1

r = 0.03 m

R = 3 + 5 = 8 cm = 0.08 m

L = length of wire = 10 cm = 0.10 m

F1 = Force on the upper side

F1 = (\mu/4\pi) (2 i2L /r) = (10-7) (2 (3.5)2 (0.10) /(0.03)) = 81.7 x 10-7 N upward

F2 = Force on the lower side

F2 = (\mu/4\pi) (2 i2L /R) = (10-7) (2 (3.5)2 (0.10) /(0.08)) = 30.625 x 10-7 N downward

Net force is given as

Fnet = F1 - F2 = 81.7 x 10-7 - 30.625 x 10-7 = 51.075 x 10-7 N = 5.1 x 10-6 N